OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5

OpenStudy (anonymous):

a) -4x - 9y = -19 -10y = -30 b)4x + 3y = 5 2y = -6 c) 7x - 3y = -11 9x = -6 d)12x - 3y = -33 14x = -28

OpenStudy (anonymous):

y =4x + 11 put this equation in 2nd eq. and solve for x then use that value of x in first eq. to get the value of y then put those values in the given equations to find the correct eq

OpenStudy (anonymous):

how do i put it in the second equation

OpenStudy (anonymous):

2x +3(4x +11) =5

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

correct. now plug this value in first eq. to find y.

OpenStudy (anonymous):

how do i do that

OpenStudy (compassionate):

It seems he has abandoned you.

OpenStudy (anonymous):

i see D:

OpenStudy (anonymous):

what part do i plug it in

OpenStudy (anonymous):

1st equation. 4x - y = -11 sorry net prob

OpenStudy (anonymous):

would it be 2x-y=-11

OpenStudy (anonymous):

4(-2) - y = -11 4(-2) + 11 = y

OpenStudy (anonymous):

3=y

OpenStudy (anonymous):

Correct. Now which eq in the given option will have this as a solution??

OpenStudy (anonymous):

@Compassionate dear, its her not him. ;)

OpenStudy (anonymous):

2y=6 would be y=3 becasue you would devide both sides by 2 to make the y alone right

OpenStudy (anonymous):

you y = 3 not -3 so no its not second

OpenStudy (anonymous):

oohhh

OpenStudy (anonymous):

Sure??

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

a) -4(-2) - 9(3) = 8 - 27 =-19 -10(3) = -30 b)4(-2) + 3(3) = -8 + 9 != 5 2(3) != -6 c) 7(-2) - 3(3) = -14 -9 =-23 9(3) = 27 d)12(-2) - 3(3) = -24 - 9 =-33 14(-2) = -28 its a and d

OpenStudy (anonymous):

its both?

OpenStudy (anonymous):

no because d would be 2 not 3

OpenStudy (anonymous):

well negative 2

OpenStudy (anonymous):

14 x = - 28 14(-2) = -28

OpenStudy (anonymous):

so its a then..

OpenStudy (anonymous):

yeah and also d is it 14x or 14y in eq d

OpenStudy (anonymous):

oh well thank you alot i am begining to get this topic now

OpenStudy (anonymous):

anytime :)